# A Step in Flux to Suppress Axion Isocurvature

### v1

@@ 1.1 The axion isocurvature constraint
${\cal P}_{{\cal R}{\cal R}}(k)=A_{s}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},\quad A_{s}\approx 2.1\times 10^{-9},\quad n_{s}\approx 0.965,$ (3)
@@
where $\gamma_{\mathrm{dil}}$ is a possible dilution factor arising from late-time entropy production, which is a reasonable expectation in supersymmetric settings from saxion or moduli decays (see, e.g., [45, 46, 47, 48, 49]), and $\gamma_{\mathrm{anh}}$ is a correction for anharmonic effects in the potential, most important if the initial $\theta_{i}$ is close to $\pi$ (see, e.g., [50, 51, 52]). For large values of $f_{a}$, the correct amount of dark matter abundance can be achieved by taking $\theta_{i}\ll 1$ (often referred to as the “anthropic axion window” [53, 54]) or $\gamma_{\mathrm{dil}}\ll 1$ (though this is limited by the requirement that the universe reheat above the BBN temperature). A third alternative, the stochastic axion scenario [55, 56], requires an extremely low $H_{I}$ and is not relevant for our goals in this paper. Combining (4), (5) and (6), we obtain the following constraints on the scenario in which axions constitute only a fraction of dark matter, and on two scenarios in which they constitute all of dark matter:
[…]
@@ 1.2 Relaxing the axion isocurvature constraint
    Time-dependent decay constant. $f_{I}\gg f_{a}$. Because the axion isocurvature fluctuations are proportional to $H_{I}/(2\pi f_{I})$, taking $f_{I}$ larger can reduce isocurvature, whereas only the late-time value $f_{a}$ determines the dark matter abundance, as is apparent in (8) and (9). This has been considered in many papers [88, 53, 89, 90, 91, 92, 93, 94, 95, 96, 81, 97, 98]. For an extra-dimensional axion, it is plausible that this scenario could be realized by the dynamics of the volume modulus [99, 100].
@@ 2.1 Review: axion monodromy mass
Figure 2: Axion monodromy potential in effective theory: $V_{\mathrm{eff}}(\theta,j)=\frac{1}{2}m_{\theta}^{2}f^{2}\left((\theta-\theta_{0})-2\pi j/n\right)^{2}$. The potential exhibits monodromy when $\theta\to\theta+2\pi/n$, $j\to j+1$. Dynamical $A$-branes interpolate between branches with different $j$.
@@
The gauge transformation $\theta\mapsto\theta+2\pi$ is accompanied by a monodromy permuting the branches, $j\mapsto j+n$ [^3] This is an instance of the standard result that Chern-Simons terms dualize to Stueckelberg terms, in a slightly non-standard form where $j$ is viewed as a field strength even though there is no obvious way to define a $(-1)$-form gauge field for it to be the field strength of. Readers can find a pedagogical review of Chern-Simons terms, Stueckelberg terms, and dualities between them in [31]. This theory can admit domain walls, electrically charged under $A_{3}$, across which the value of $j$ jumps. We refer to these domain walls as $A$-branes. A non-canonical kinetic term for $A_{3}$ would lead to a non-quadratic effective potential for $\theta$; while this may seem to go beyond the validity of EFT, these terms may be determined by some organizing principle like the DBI action that nonlinearly realizes higher-dimensional Poincaré invariance [104]. Useful earlier references related to four-form gauge field strengths, associated domain walls, and effects on axion mass include [107, 108, 109, 110, 111].
@@ 2.2 Making the axion monodromy mass dynamical
$-\int\frac{1}{2e_{B}^{2}}\left[\textrm{d}B_{3}-(\theta-\theta_{0})F_{4}\right]\wedge\star\left[\textrm{d}B_{3}-(\theta-\theta_{0})F_{4}\right].$ (15)
@@
$\widetilde{H}_{4}\equiv\textrm{d}B_{3}-(\theta-\theta_{0})F_{4}.$ (17)
@@ 3.1 Bubble nucleation and percolation
Figure 4: The process of $B$-brane bubble mergers. As explained in [119], when bubbles merge, the overlap region will have a different flux ($n=-1$ in this example). Left panel: the state immediately after bubbles collide. Arrows indicate the direction of pressure on the bubble walls due to vacuum energy differences, assuming that the $n=0$ state has lowest vacuum energy. Right panel: we expect that the branes will reconnect and interior regions with $n\neq 0$ will collapse, eventually leading to a universe in the $n=0$ state everywhere.
@@
An important physical effect in flux tunneling, emphasized in [119], is that because the value of $n$ changes across a $B$-brane, the merger of bubbles creates an intermediate state with a different value of $n$, as illustrated in the left panel of Fig. 4. If the phase transition is between an $n=1$ state and an $n=0$ state, the merger of two bubbles produces an intermediate bubble of the $n=-1$ state. We assume, crucially, that the $n=0$ state (in which we live today) has lower energy than all $n\neq 0$ states. [^7] Much of the literature on this topic assumes Freund-Rubin-like compactifications, in which the flux being studied is balanced against curvature to stabilize the extra dimensions. Such cases are incompatible with our assumption of a minimum (and finite) energy $n=0$ state. We have in mind a UV completion where the flux $n$ is through a small cycle within a larger compactification and has little effect on the overall volume, so that we do not have to make the same assumptions about the $n$-dependence of $V_{n}$ made in other works. It is then energetically favored for the bubble walls to reconnect, allowing the $n=-1$ region to pinch off and collapse, as in the right panel of Fig. 4. As it collapses, it will radiate away energy into light fields. If the phase transition occurs at or after the end of inflation, the collapse of such negative-$n$ bubbles could play an important role in reheating, though it is difficult to make a more precise statement without a specific UV completion. The collision of bubble walls could also produce primordial gravitational waves, which may provide a detectable signal of a first-order phase transition. We will briefly comment on this in our concluding section 7.
@@ 4 Case Study: Tunneling During Inflation
Let us first establish some generalities. We will be dealing with a scalar operator $\hat{X}_{\mathbf{k}}(\tau)$ with co-moving momentum $\mathbf{k}$, where $\hat{X}\in\{\delta\hat{\theta},\delta\hat{\phi}\}$ is either the axion or the inflaton fluctuations, and assume that the state for each is its standard Bunch-Davies vacuum. The operator can be written in terms of its mode functions
@@ 5 Constraints for $(H_{I},f_{a})$
The red line shows the upper bound on $H_{I}$ obtained from $r$ (10). The observation of a neutrino pulse from SN1987a places a lower bound of $f_{a}\gtrsim 4\times 10^{8}\leavevmode\nobreak\ \mathrm{GeV}$ [141, 142], which is plotted in orange. We also require $f_{a}<M_{\mathrm{Pl}}$, as shown in the teal line. This is an expected constraint on valid quantum gravity theories [143]. [^10] One actually expects a stronger bound, $f_{a}\lesssim M_{\mathrm{Pl}}/S_{\mathrm{inst}}$ [144] (the axion Weak Gravity Conjecture), but the precise numerical coefficient is somewhat uncertain [145] and may depend on the details of mixing of the QCD axion with other axions [114]. For the fractional DM scenario, the yellow curve represents the constraint $\Omega_{a}h^{2}<\Omega_{c}h^{2}=0.12$. The axion dark matter abundance is given by Eq. (6), where
@@
${\cal P}_{{\cal I}{\cal I}}(k)=A_{I}\left(\frac{k}{k_{*}}\right)^{n_{I}-1},$ (63)
and is obtained at $k_{*}=0.05\text{ Mpc}^{-1}$. Our mechanism predicts a different power spectrum (57) which critically depends on $k_{*}$. Since this is the same as the broken power law case studied in [146], we can use their limits on $(\Omega_{a}/\Omega_{c})^{2}A_{I}$ to bound $\beta_{\text{iso}}(k_{*})$. Here we choose several different values of $k_{*}$ and plot the curves that saturate $\beta_{\text{iso}}(k_{*})$ in green, with different line styles.
@@ 6.1 Dynamical integer from extra-dimensional flux
We begin with a Chern-Simons term in a $d$-dimensional theory, with the $(d-4)$ extra dimensions compactified on a manifold $Y$. This theory contains a Chern-Simons term,
@@
with $p+q+r=d-2$ and $\ell\in{\mathbb{Z}}$. (It is possible that some of the fields $A$, $\widetilde{B}$, and $C$ are the same, which would change the normalization of the prefactor.) The 4d fields are related to the $d$-dimensional fields via:
@@
  • $A_{3}$ is a 4d zero mode from reducing $A_{r}$ along an $(r-3)$-cycle $\Gamma_{r-3}$,
@@
The object magnetically charged under $\widetilde{B}_{q}$ is a $(d-q-3)$-brane, or equivalently a $(p+r-1)$-brane. If we wrap this on the $p$-cycle and the $(r-3)$-cycle we are left with a 2-brane in 4 dimensions, i.e., a domain wall. This is the object across which the flux of $\textrm{d}\widetilde{B}_{q}$ jumps, and accordingly the integer $n$ changes. It is electrically charged under the dual $B$ of $\widetilde{B}$, which is a 3-form gauge field in four dimensions, as in (15). This object is the $B$-brane that must be nucleated to produce our cosmological phase transition.

### v2

@@ 1.1 The axion isocurvature constraint
${\cal P}_{{\cal R}{\cal R}}(k)=A_{s}\left(\frac{k}{k_{\text{pivot}}}\right)^{n_{s}-1},\quad A_{s}\approx 2.1\times 10^{-9},\quad n_{s}\approx 0.965,$ (3)
@@
where $\gamma_{\mathrm{dil}}$ is a possible dilution factor arising from late-time entropy production, which is a reasonable expectation in supersymmetric settings from saxion or moduli decays (see, e.g., [165, 121, 122, 104, 19, 85, 17]), and $\gamma_{\mathrm{anh}}$ is a correction for anharmonic effects in the potential, most important if the initial $\theta_{i}$ is close to $\pi$ (see, e.g., [169, 132, 171]). For large values of $f_{a}$, the correct amount of dark matter abundance can be achieved by taking $\theta_{i}\ll 1$ (often referred to as the “anthropic axion window” [127, 87]) or $\gamma_{\mathrm{dil}}\ll 1$ (though this is limited by the requirement that the universe reheat above the BBN temperature). A third alternative, the stochastic axion scenario [80, 168], requires an extremely low $H_{I}$ and is not relevant for our goals in this paper. Combining (4), (5) and (6), we obtain the following constraints on the scenario in which axions constitute only a fraction of dark matter, and on two scenarios in which they constitute all of dark matter:
$\text{Fractional DM:}\quad H_{I}$ $<\left(\frac{f_{I}}{\theta_{i}f_{a}}\right)\,\left(\frac{f_{a}}{10^{12}\,\mathrm{GeV}}\right)^{-0.165}\,2.5\times 10^{7}\,\mathrm{GeV},$ (7)
$\text{Anthropic:}\quad H_{I}$ $<\left(\frac{f_{I}}{f_{a}}\right)\,\left(\frac{f_{a}}{10^{12}\,\mathrm{GeV}}\right)^{0.4175}\,2.6\times 10^{7}\,\mathrm{GeV},$ (8)
$\text{Dilution:}\quad H_{I}$ $<\theta_{i}\,\left(\frac{f_{I}}{10^{12}\,\mathrm{GeV}}\right)\,2.8\times 10^{7}\,\mathrm{GeV}.$ (9)
@@ 1.2 Relaxing the axion isocurvature constraint
    Time-dependent decay constant. $f_{I}\gg f_{a}$. Because the axion isocurvature fluctuations are proportional to $H_{I}/(2\pi f_{I})$, taking $f_{I}$ larger can reduce isocurvature, whereas only the late-time value $f_{a}$ determines the dark matter abundance, as is apparent in (8) and (9). This has been considered in many papers [124, 127, 100, 101, 72, 112, 50, 70, 143, 83, 113, 116, 7, 79]. For an extra-dimensional axion, it is plausible that this scenario could be realized by the dynamics of the volume modulus [59, 44].
@@ 2.1 Review: axion monodromy mass
Figure 2: Axion monodromy potential in effective theory: $V_{\mathrm{eff}}(\theta,j)=\frac{1}{2}m_{\theta}^{2}f^{2}\left((\theta-\theta_{0})-2\pi j/n\right)^{2}$. The potential exhibits monodromy when $\theta\to\theta+2\pi/n$, $j\to j+1$. Dynamical AA-branes interpolate between branches with different $j$.
@@
The gauge transformation $\theta\mapsto\theta+2\pi$ is accompanied by a monodromy permuting the branches, $j\mapsto j+n$. [^3] This is an instance of the standard result that Chern-Simons terms dualize to Stueckelberg terms, in a slightly non-standard form where $j$ is viewed as a field strength even though there is no obvious way to define a (1)(-1)-form gauge field for it to be the field strength of. Readers can find a pedagogical review of Chern-Simons terms, Stueckelberg terms, and dualities between them in [154]. This theory can admit domain walls, electrically charged under $A_{3}$, across which the value of $j$ jumps. We refer to these domain walls as $A$-branes. A non-canonical kinetic term for $A_{3}$ would lead to a non-quadratic effective potential for $\theta$; while this may seem to go beyond the validity of EFT, these terms may be determined by some organizing principle like the DBI action that nonlinearly realizes higher-dimensional Poincaré invariance [163]. Useful earlier references related to four-form gauge field strengths, associated domain walls, and effects on axion mass include [33, 34, 75, 30, 68].
@@ 2.2 Making the axion monodromy mass dynamical
$\int\frac{1}{2e_{B}^{2}}\left[\textrm{d}B_{3}-\frac{1}{2\pi}(\theta-\theta_{0})F_{4}\right]\wedge\star\left[\textrm{d}B_{3}-\frac{1}{2\pi}(\theta-\theta_{0})F_{4}\right].$ (15)
@@
$\widetilde{H}_{4}\equiv\textrm{d}B_{3}-\frac{1}{2\pi}(\theta-\theta_{0})F_{4}.$ (17)
@@ 3.1 Bubble nucleation and percolation
Figure 4: The process of $B$-brane bubble mergers. As explained in [28], when bubbles merge, the overlap region will have a different flux (n=1n=-1 in this example). Left panel: the state immediately after bubbles collide. Arrows indicate the direction of pressure on the bubble walls due to vacuum energy differences, assuming that the $n=0$ state has lowest vacuum energy. Right panel: we expect that the branes will reconnect and interior regions with $n\neq 0$ will collapse, eventually leading to a universe in the $n=0$ state everywhere.
@@
An important physical effect in flux tunneling, emphasized in [28], is that because the value of $n$ changes across a $B$-brane, the merger of bubbles creates an intermediate state with a different value of $n$, as illustrated in the left panel of Fig. 4. If the phase transition is between an $n=1$ state and an $n=0$ state, the merger of two bubbles produces an intermediate bubble of the n=1n=-1 state. We assume, crucially, that the $n=0$ state (in which we live today) has lower energy than all $n\neq 0$ states. [^7] Much of the literature on this topic assumes Freund-Rubin-like compactifications, in which the flux being studied is balanced against curvature to stabilize the extra dimensions. Such cases are incompatible with our assumption of a minimum (and finite) energy $n=0$ state. We have in mind a UV completion where the flux $n$ is through a small cycle within a larger compactification and has little effect on the overall volume, so that we do not have to make the same assumptions about the $n$-dependence of $V_{n}$ made in other works. It is then energetically favored for the bubble walls to reconnect, allowing the n=1n=-1 region to pinch off and collapse, as in the right panel of Fig. 4. As it collapses, it will radiate away energy into light fields. If the phase transition occurs at or after the end of inflation, the collapse of such negative-$n$ bubbles could play an important role in reheating, though it is difficult to make a more precise statement without a specific UV completion. The collision of bubble walls could also produce primordial gravitational waves, which may provide a detectable signal of a first-order phase transition. We will briefly comment on this in our concluding section 7.
@@ 4 Case Study: Tunneling During Inflation
[…]
@@ 5 Constraints for $(H_{I},f_{a})$
The red line shows the upper bound on $H_{I}$ obtained from $r$ (10). The observation of a neutrino pulse from SN1987a places a lower bound of $f_{a}\gtrsim 4\times 10^{8}\penalty 10000\ \mathrm{GeV}$ [151, 42], which is plotted in orange. We also require $f_{a}<M_{\mathrm{Pl}}$, as shown in the teal line. This is an expected constraint on valid quantum gravity theories [18]. [^10] One actually expects a stronger bound, $f_{a}\lesssim M_{\mathrm{Pl}}/S_{\mathrm{inst}}$ [12] (the axion Weak Gravity Conjecture), but the precise numerical coefficient is somewhat uncertain [84] and may depend on the details of mixing of the QCD axion with other axions [27]. For the fractional DM scenario, the yellow curve represents the constraint $\Omega_{a}h^{2}<\Omega_{c}h^{2}=0.12$. The axion dark matter abundance is given by Eq. (6), where
@@
${\cal P}_{{\cal I}{\cal I}}(k)=A_{I}\left(\frac{k}{k_{\text{pivot}}}\right)^{n_{I}-1},$ (63)
and is obtained at $k_{\text{pivot}}=0.05\text{ Mpc}^{-1}$. Our mechanism predicts a different power spectrum (57) which critically depends on $k_{*}$. Since this is the same as the broken power law case studied in [37], we can use their limits on $(\Omega_{a}/\Omega_{c})^{2}A_{I}$ to bound $\beta_{\text{iso}}(k_{*})$. Here we choose several different values of $k_{*}$ and plot the curves that saturate $\beta_{\text{iso}}(k_{*})$ in green, with different line styles.
@@ 6.1 Dynamical integer from extra-dimensional flux
We begin with a Chern-Simons term in a $d$-dimensional theory, with the (d4)(d-4) extra dimensions compactified on a manifold $Y$. This theory contains a Chern-Simons term,
@@
with p+q+r=d2p+q+r=d-2 and $\ell\in{\mathbb{Z}}$. (It is possible that some of the fields $A$, $\widetilde{B}$, and $C$ are the same, which would change the normalization of the prefactor.) The 4d fields are related to the $d$-dimensional fields via:
@@
  • $A_{3}$ is a 4d zero mode from reducing $A_{r}$ along an (r3)(r-3)-cycle $\Gamma_{r-3}$,
@@
The object magnetically charged under $\widetilde{B}_{q}$ is a (dq3)(d-q-3)-brane, or equivalently a (p+r1)(p+r-1)-brane. If we wrap this on the $p$-cycle and the (r3)(r-3)-cycle we are left with a 2-brane in 4 dimensions, i.e., a domain wall. This is the object across which the flux of $\textrm{d}\widetilde{B}_{q}$ jumps, and accordingly the integer $n$ changes. It is electrically charged under the dual $B$ of $\widetilde{B}$, which is a 3-form gauge field in four dimensions, as in (15). This object is the $B$-brane that must be nucleated to produce our cosmological phase transition.